OpenStudy (anonymous):

Simplify the expression of 5x-6y+4x+3y=

OpenStudy (anonymous):

5x + 4x -6y + 3y =

OpenStudy (anonymous):

simply add/subtract the terms with the same variable (x or y)

jigglypuff314 (jigglypuff314):

5x + 4x = 9x -6y + 3y = 3y